1) A recent study of 750 internet users in Europe found that 260 of internet users were women. What is the 95% confidence interval of the true proportion of women in Europe who use the internet?
2) A survey of 800 women shoppers found that 17% of them shop on impulse. What is the 98& confidence interval for the true proportion of women shhoppers who shop on impulse?

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

1) A recent study of 750 internet users in Europe found that 260 of internet users were women. What is the 95% confidence interval of the true proportion of women in Europe who use the internet?
----------------------
sample proportion: 260/750 = 0.3467
ME = 1.96*sqrt[0.3467*0.6533/750] = 0.0744
---
95% CI: 0.35-0.07 < p < 0.35+0.07
=============================================

2) A survey of 800 women shoppers found that 17% of them shop on impulse. What is the 98& confidence interval for the true proportion of women shoppers who shop on impulse?
---
Same procedure as above.
==================
Cheers,
Stan H.
=================

Answer by edjones(8007)

n=750, p=260/750=.347, q=.653
(a,b)=p+z[alpha/2]*sqrt((p*q)/n)
=.347+-(1.96)*sqrt((.347*.653)/750) Z for the 95% confidence level is 1.96
=.347+-(1.96)(.0174)
=.347+-.034
=(.381, .313)
.
Ed